OpenStudy (anonymous):

simplify (32-27)^3+(19+1)^3 please show work i have no idea how to do this

OpenStudy (bibby):

PEDMAS. simplify what's in the parentheses first

OpenStudy (gorv):

subtract 27 from 33 and multiply it 3 times with itself and add 19 and 1 and multiply it 3 times

rishavraj (rishavraj):

a^3 + b^3 = (a +b)*(a^2 + b^2 - ab)

OpenStudy (mikkel900):

Hello, Miss. I am not entirely sure, but I think you would do like this: 32 - 27 = 5 19 + 1 = 20 So you get this: 5^3 + 20^3 And do this: 5^3 = 125 (5*5=25, 25*5=125) 20^3 = 8000 (Same method as above) Now it looks like this: 125 + 8000 You add them together, and you get 8125 Once again, I am not 100 % sure, but that is the way I would do it. Good luck!